### v1

# IBIS: A Powerful Hybrid Architecture for Human Activity Recognition

Abstract
The increasing interest in Wi-Fi sensing stems from its potential to capture environmental data in a low-cost, non-intrusive way, making it ideal for applications like healthcare, space occupancy analysis, and gesture-based IoT control. However, a major limitation in this field is the common problem of overfitting, where models perform well on training data but fail to generalize to new data. To overcome this, we introduce a novel hybrid architecture that integrates Inception-BiLSTM with a Support Vector Machine (SVM), which we refer to as IBIS. Our IBIS approach is uniquely engineered to improve model generalization and create more robust classification boundaries. By applying this method to Doppler-derived data, we achieve a movement recognition accuracy of nearly 99%. Comprehensive performance metrics and confusion matrices confirm the significant effectiveness of our proposed solution.
I Introduction
Wi-Fi sensing is an increasingly important technology that leverages ubiquitous Wi-Fi signals for non-intrusive human activity Recognition (HAR) [1]. Unlike traditional methods that rely on cameras or wearable sensors, Wi-Fi sensing utilizes Channel State Information (CSI) to capture and analyze subtle changes in signal propagation caused by human movement [2]. This approach is not only cost-effective but also preserves user privacy, making it highly suitable for applications in healthcare, surveillance, and smart homes.
However, a key challenge in this field is overfitting, where models trained on a specific dataset fail to generalize to new environments or unseen data. Many existing solutions, particularly those based on Convolutional Neural Networks (CNNs) [3] [4], are effective at capturing local patterns but often struggle to learn the broader, temporal dependencies critical for robust generalization. Furthermore, raw CSI data frequently contains noise from hardware artifacts, such as phase and frequency oscillations, which can significantly hinder accurate movement interpretation and requires a rigorous sanitization process.
To address these limitations, we propose a novel hybrid neural network architecture named Inception-BiLSTM (IBIS). The Inception component efficiently extracts local features from the input data [5], while the Bidirectional Long Short-Term Memory (BiLSTM) component captures long-range temporal dependencies. To further enhance the model’s predictive power and refine classification boundaries, we integrate a SVM with a non-linear kernel in a post-processing stage. This ensemble learning approach [6] leverages the strengths of multiple algorithms to achieve a more robust and reliable solution.
Our model’s effectiveness was evaluated on five and eight human activity classification tasks using Doppler-derived data. We compared our IBIS approach against two state-of-the-art methods: the SHARP algorithm [7], a recent ensemble-based method for HAR, and an Attention-Based BiLSTM (ABLSTM) approach [8] that uses raw CSI. The results demonstrate that our hybrid Inception-BiLSTM network, augmented with a non-linear SVM, significantly outperforms both benchmark models. Our method achieved an average accuracy of nearly 99%, validating the benefits of combining Doppler features with a sophisticated hybrid deep learning architecture for enhanced HAR performance, particularly in challenging generalization scenarios.
The remainder of this paper is structured as follows. Section II presents a comprehensive overview of existing research on HAR using hybrid neural networks, providing a fresh perspective on the field. Section III details the development and implementation of our proposed algorithms, including the specific neural network architecture. The validation process, which includes activity classification confusion matrices and Receiver Operating Characteristic (ROC) curve analysis, is described in Section IV. Finally, Section V provides a summary of our findings and offers concluding remarks.
II Related Works
This section provides a review of state-of-the-art research in Wi-Fi sensing, focusing on the machine learning algorithms commonly used for CSI data analysis. We also discuss ensemble learning techniques that have been applied to optimize model performance. Recent studies have explored various approaches to enhance Wi-Fi sensing capabilities. Table I presents a comparative overview of the models and key features from previous studies on CSI-based Human Activity Recognition.
Son et al. [9] improved indoor occupancy detection by proposing a preprocessing method that uses nonlinear transformations to expand CSI data into three channel vectors, which amplified signal differences. A CNN was then applied, achieving an average accuracy of approximately 95%. This work highlights the importance of effective signal preprocessing for boosting classification performance.
Ensemble learning has emerged as a key strategy to improve the reliability and accuracy of machine learning models. Wei et al. [10] used an ensemble learning approach for HAR by combining feature extraction with an SVM classifier. This method reportedly outperformed other literature approaches, reaching an accuracy of about 98.95%.
Similarly, Ye et al. [11] developed a system for intrusion monitoring that combines Deep Neural Networks (DNNs) with BiLSTM networks. While their approach demonstrated improved performance over traditional methods, they noted a significant weakness in detecting specific actions, such as falling movements, with an accuracy of only 62.65%.
The SHARP algorithm, as proposed by Meneghello et al. [7], provides a foundational method for HAR based on Doppler traces. Their work addressed noise and hardware interference and, after applying a neural network-based refinement, achieved an accuracy of up to 95%. The authors also demonstrated that antenna merging could further enhance accuracy. Building upon this, Cominelli et al. [12] investigated the impact of Wi-Fi 6 features and developed a large-scale dataset for benchmarking. However, they reported difficulties in reproducing the SHARP algorithm’s results, suggesting a sensitivity to environmental variability. Their findings also indicated that SHARP’s performance degrades with increased bandwidth and a higher number of activity classes.
Despite these advancements, a major challenge in Wi-Fi sensing remains the issue of poor generalization to unseen data. He et al. [8] addressed this limitation by introducing a cross-subject transfer learning method based on a CNN-ABLSTM model. By training on separate source and target domains, they achieved an accuracy of 85.82% in cross-subject recognition tasks, demonstrating a promising path for creating more robust and generalizable models.
TABLE I: Comparison of Models and Features in Human Activity Recognition Using CSI
Specification [9] [10] [7] [12] [11] [8]
Model CNN CNN+RF+SVM Inception Inception DNN-ABLSTM CNN-ABLSTM
Accuracy 95% 98.85% 88.12% 77.72% 92.06% 85.82%
Application Indoor Occupancy HAR HAR HAR HAR HAR
Network 802.11 802.11n 802.11ac 802.11ac/ax 802.11n -
Bandwidth 20MHz - 80MHz 80/160MHz - -
CSI Data Amplitude Temporal and Spatial Correlation Phase CFR Phase CFR Amplitude and Phase Raw CSI
III System Modeling
III-A Dataset
The dataset for this research is based on the SHARP study [7] and consists of CSI data collected from three volunteers (one male, two females) performing eight activities in various indoor environments. These activities: empty, sitting, stand up, walking, running, jumping, and arm gym—were each recorded for 120 seconds. The data was gathered in a bedroom, living room, and university laboratory during two periods: April–December 2020 and January 2022. A Nexmon CSI Extraction Tool and an Asus RT-AC86U Wi-Fi router were used to extract the CSI from Wi-Fi 2.4 GHz preamble headers.
CSI is a data framework that describes how a wireless signal travels from a transmitter to a receiver. It provides high-resolution data on the physical propagation properties of a signal, including the effects of diffraction, reflection, and scattering [9]. Because CSI captures both amplitude and phase information, it is sensitive to environmental changes, such as the movement of people.
The dataset includes seven distinct scenarios (S1-S7):
  • S1–S3: A bedroom with a central bookcase. S1 was used for model training, S2 for a generalization scenario, and S3 was measured on different days than S1;
  • S4–S5: Activities performed alternately within the bedroom;
  • S6: Data collected in a living room on different days;
  • S7: A laboratory with numerous desks, computers, and monitors, measured over several days.
Scenario S7 was chosen as the benchmark for this study’s model evaluation because it presented the greatest challenge. According to the SHARP authors, the high density of obstacles and reflective surfaces in this environment significantly impacts CSI signal quality and classification performance.
III-B Decision-Making Process and Algorithm Steps
To understand the purpose of our work, it is essential to first comprehend the original SHARP method. The SHARP process begins with the collection of CSI data from three individuals, with each of the eight activities sampled for 120 seconds. After the CSI data is extracted from the Wi-Fi 2.4 GHz packet preambles, it undergoes three key stages, as shown in Figure  1:
  1. 1.
    Pre-processing: This stage sanitizes the CSI data by eliminating noise, interference, and phase ambiguity. A reference signal path is used to remove phase offsets, recovering a clean signal that reveals the Doppler traces associated with movement;
  2. 2.
    Classification: The sanitized Doppler traces are fed into an Inception neural network. To improve generalization beyond the original architecture’s spatial feature learning, the network was enhanced with a BiLSTM layer, creating a hybrid Inception-BiLSTM model. This modification allows the model to capture both spatial and temporal dependencies, significantly boosting performance in human activity recognition;
  3. 3.
    Post-processing: The original SHARP method used a majority voting mechanism among the four receiving antennas to determine the final activity label. This step smoothed outliers and improved robustness.
In this work, we modified the post-processing phase by introducing a Support Vector Machine (SVM) before the antenna fusion stage. This ensemble learning strategy leverages the SVM to enhance classification boundaries, thereby improving the overall precision and reliability of the final decision.
Figure 1: The sequential process diagram illustrates the IBIS steps: collection of raw CSI data, passing through sanitization, neural network training, and Ensemble Learning post-Processing.
III-C Hybrid Inception-BiLSTM (IBIS) Network
To interpret and classify data from Doppler traces, we designed a hybrid neural architecture combining two networks. The Inception network processes global features and learns spatial characteristics, while the BiLSTM network captures temporal dependencies and patterns. This combination is designed to optimize both classification performance and generalization ability.
Figure 2 illustrates the integrated architecture of these two neural structures. While some layers like Normalization, Reshape, and Dropout are omitted for visualization, they are described in the text. Table II provides the input and output dimensions and parameter counts for each block. The shaded and hatched purple areas in the figure highlight the novel components of our proposal. Our approach enhances the standard Inception network, which typically only includes activation and convolution layers for learning local features. We added a MaxPooling layer and Batch Normalization to improve convergence and filter relevant features. Furthermore, a BiLSTM network was incorporated to enable the model to learn temporal patterns and motion dynamics from Doppler shifts.
III-C1 Inception Network
The Inception network is configured to extract spatial features from the Doppler trace data. A description of each block and its function is as follows:
  • Input: The entry point for the Doppler trace data;
  • Normalization: Normalizes input data by subtracting the mean and dividing by the standard deviation to stabilize training and improve convergence speed;
  • Conv2D (Conv Block): This block combines Conv2D, Batch Normalization, and an Activation function. The Conv2D layer uses filters to extract local patterns like edges and textures;
  • Batch Normalization: Normalizes activations to accelerate training and regularize the model;
  • Activation: Introduces non-linearity, allowing the model to learn complex representations;
  • Max Pooling: Reduces spatial resolution and complexity to focus on the most relevant features;
  • Concatenate: Merges activation maps from different paths, maintaining important information while reducing dimensionality;
  • Dropout: Randomly deactivates neurons during training to reduce overfitting and improve generalization;
  • Reshape: Transforms the output tensor into a sequence format, preparing it for temporal processing by the BiLSTM layer.
The combination of Conv2D, Batch Normalization, and Activation is repeated nine times within the Inception module, a core characteristic that increases depth and improves generalization by learning more abstract features.
III-C2 BiLSTM Network
The second part of the architecture is the BiLSTM network, which was included to capture crucial temporal patterns and dependencies from the Doppler traces. These traces are time-series signals that reflect movement and channel variations, often exhibiting complex patterns that the network must recognize.
The BiLSTM network’s components are described as follows:
  • BiLSTM + Attention: This layer processes input sequences in both forward and backward directions, capturing complex temporal relationships. The attention mechanism enhances the model’s focus on the most relevant time steps;
  • Global Average Pooling: Reduces the dimensionality of the feature space in preparation for the final classification layer;
  • Dense Layer: A fully connected layer that generates the final output, with the number of units depending on the number of output classes (Table II).
While the CNN-ABLSTM model from a comparative study [8] has a lower computational cost due to its lighter architecture, it only achieves an accuracy of 85.82%, compared to the 98.95% of our IBIS method. This performance gap is attributed to the limited number of attention layers in their model, which are insufficient to capture the complex temporal and spatial patterns and subtle variations present in data with high variability and Doppler effects. Our hybrid Inception-BiLSTM network, with its greater depth and robust attention mechanisms, is better equipped to handle complex local features, leading to higher learning capacity and a more efficient convergence rate. This is because the Inception module learns spatial features, while the BiLSTM module captures temporal dependencies, making the combined model more robust to noise and interference, which is ideal for classifying Doppler traces.
Figure 2: Architecture of the IBIS hybrid neural network for Human Activity Recognition using Doppler traces, showing all convolutional and attention layers.
TABLE II: Inception-BiLSTM Architecture (Compact Notation)
Block / Layer Output Shape Parameters
Input Layer (32, 32, 3) -
Normalization (32, 32, 3) mean, variance
Conv2D(3×1×3) (32, 32, 3) 3
BatchNormalization (32, 32, 3) 12
Activation (swish) (32, 32, 3) -
Conv2D(6×2×3) (31, 31, 6) 78
BatchNormalization (31, 31, 6) 24
Activation (swish) (31, 31, 6) -
Conv2D(5×2×6) (30, 30, 5) 125
BatchNormalization (30, 30, 5) 20
Conv2D(9×4×6) (28, 28, 9) 873
BatchNormalization (28, 28, 9) 36
Activation (swish) (30×30×5) / (28×28×9) -
MaxPooling2D(2×2) (15, 15, 5) -
Concatenate (15, 15, 14) -
Conv2D(3×1×14) (15, 15, 3) 45
BatchNormalization (15, 15, 3) 12
Activation (swish) (15, 15, 3) -
Conv2D(6×2×3) (14, 14, 6) 78
BatchNormalization (14, 14, 6) 24
Conv2D(5×2×6) (13, 13, 5) 125
BatchNormalization (13, 13, 5) 20
Conv2D(9×4×6) (11, 11, 9) 873
BatchNormalization (11, 11, 9) 36
Activation (swish) (13×13×5) / (11×11×9) -
MaxPooling2D(2×2) (6, 6, 5) -
Concatenate (6, 6, 14) -
Dropout(0.5) (6, 6, 14) -
Conv2D(6×1×14) (6, 6, 6) 84
BatchNormalization (6, 6, 6) 24
Activation (swish) (6, 6, 6) -
Reshape (36, 6) -
Bidirectional LSTM (64 units) (36, 128) 45568
Activation (tanh) (36, 128) -
Attention (36, 128) 16640
GlobalAvgPooling1D (RNN) (128,) -
GlobalAvgPooling1D (Attention) (128,) -
Concatenate (256,) -
Dropout(0.5) (256,) -
Dense (5 classes) (5,) 1285
Note: ‘Conv2D(out×kernel×in)‘ notation refers to output channels, filter size (square or rectangular kernel), and input channels, respectively. Shapes may vary slightly depending on padding/stride.
III-D Support Vector Machine with Radial Basis Function Kernel
The hybrid IBIS network is effective at detecting and classifying both global patterns and temporal sequences. However, for accurate classification of CSI-derived data, this architecture alone is insufficient. To address this, a SVM was integrated into the post-processing stage as an ensemble learning strategy. Unlike common approaches where SVM is the primary classifier, this work uses it to refine the final classification results and improve generalization. The SVM’s ability to handle outliers and its use of decision margins contribute to better model generalization, which is particularly useful for Doppler trace data that is prone to noise and distortions. This allows the SVM to refine the final classification results, especially for samples that might be misclassified by the neural network.
A key advantage of SVM is its kernel function flexibility. The Radial Basis Function (RBF) kernel [13] [14] was chosen for this study’s Doppler data due to its ability to adapt decision boundaries to non-linear distributions. The penalty parameter C was set to 1, providing an optimal balance between training accuracy and generalization. Additionally, class balancing was applied to the SVM configuration to ensure that each class was given equal weight, which prevents the model from being biased toward the majority class and leads to a fairer, more robust classification. Figure 3 illustrates the SVM model’s application, showing five distinct classes separated by well-defined boundaries. Proper class segmentation is essential for understanding the unique features of each class, leading to improved classification accuracy.
Figure 3: Representation of the SVM algorithm applied to five movements with non-linear Data, applying the PCA method for dimensionality reduction.
IV Performance Evaluation
This section evaluates the hybrid IBIS neural network using generalization data from Scenario S7. This scenario, a laboratory room filled with desks, chairs, and monitors, was chosen as a benchmark due to its challenging environment. The numerous obstacles in this setup introduce complex signal distortions like multipath propagation, reflection, and scattering, which make the task of activity recognition particularly difficult for any classifier. We will analyze the model’s performance through its learning curves, class-wise confusion matrices, and the computational cost of the hybrid architecture.
IV-A Confusion Matrix for 5 Movements
Figure 4 compares the confusion matrices for five movements, showing the results from the Inception-only network against the proposed hybrid scheme. As seen in Figure 4a, the Inception network struggled with precision, particularly for the walking movement, which was frequently confused with running, achieving only 71% accuracy. Jumping was also often misclassified as running, with an accuracy of just 67%. This lack of precision is attributed to the similarity of Doppler patterns between these movements and the fact that the Inception network was not designed to handle temporal dependencies crucial for accurate CSI data classification.
Figure 4: Comparative analysis of the confusion matrix for five movements: original Inception versus the hybrid IBIS with SVM post-processing.
Conversely, Figure 4b shows the results after implementing our hybrid IBIS model and the SVM algorithm, which led to significantly better performance with only minor misclassifications. The improvement is due to the BiLSTM component’s ability to extract temporal patterns from the Doppler data. Furthermore, the SVM, with its non-linear kernel, proper C parameter tuning, and class balancing, refined the model’s classifications, leading to enhanced overall performance. Both confusion matrices presented refer to the post-antenna fusion phase, where classification decisions are combined via majority voting.
IV-B ROC Curve for 5 Movements
Figure 5 displays the learning ROC curve for all five classes, with a dotted diagonal line indicating the classification threshold. This curve was generated using generalization test data without antenna fusion, allowing for a direct evaluation of the algorithm’s generalization capability.
Figure 5: ROC curve analysis for five movements: comparing the original Inception with the hybrid Neural Network IBIS with SVM Post-Processing.
As shown in Figure 5a, the original model performed poorly on the walking, running, and jumping classes, which showed significant fluctuations in accuracy. In contrast, Figure 5b demonstrates that with our proposed approach, accuracy values improved and stabilized with no apparent oscillations, even without antenna fusion. This highlights the effectiveness of applying ensemble learning in this challenging scenario.
IV-C Confusion Matrix for 8 Movements
Figure 6 shows the confusion matrices for eight movements, including: standing, stand up, and arm gym. In Figure 6a, the original algorithm’s performance worsened compared to the five-movement scenario, with a higher misclassification rate for gestures like arm gym, jumping, and sitting. This is likely due to the SHARP algorithm’s sanitization process, where the probability of error increases proportionally with the number of classified gestures and bandwidth [12]. This behavior indicates that the original model struggled with overfitting and failed to generalize when more movements were added.
Figure 6: Comparative analysis of the confusion matrix for eight movements: original Inception versus the hybrid IBIS with SVM post-processing.
In contrast, Figure 6b demonstrates that the overall accuracy remained high despite the addition of new movements. This superior performance is attributed to the combination of the IBIS network and the SVM algorithm. Even after attempting to improve the original model’s performance by modifying the dataset split to 80% for training, 10% for testing, and 10% for validation, the results remained unchanged.
IV-D ROC Curve for 8 Movements
Figure 7 presents the ROC curves for all classes in the eight-movement scenario. The original simulation, shown in Figure 7a, performed poorly on the running and jumping classes, which exhibited constant oscillations and lower accuracy. This behavior is attributed to the model’s inability to generalize effectively, resulting in misclassifications and a higher incidence of false positives.
Figure 7: ROC curve analysis for eight movements: comparing the original Inception with the hybrid IBIS approach with SVM Post-Processing.
On the other hand, Figure 7b demonstrates significantly improved results with our proposed method. The curves show no apparent oscillations and maximum accuracy was achieved across all classes. These results confirm that combining the Inception-BiLSTM architecture with the SVM algorithm is highly effective, as the performance was considerably enhanced even without antenna fusion.
IV-E Comparative Algorithms
Table III demonstrates that the hybrid Inception-BiLSTM + SVM model achieved the best and most consistent results compared to other ensemble learning models. Its superior performance is attributed to the neural network’s ability to capture complex patterns and the SVM’s capacity to handle temporal sequence boundaries, both of which are essential for Doppler-based data. The original Inception model, for example, struggled with classifying similar movements like walking and jumping, leading to false positives and poor generalization.
TABLE III: Models Accuracy
Models Empty(%) Sitting(%) Walking(%) Running(%) Jumping(%)
IBIS+SVM 100 100 98.69 97.41 97.53
Inception 100 100 70.79 100 66.67
Note: SVM - Support Vector Machine
Similarly, Table IV, which presents average metrics for five movement classes, confirms that our IBIS + SVM model outperformed other tested models. We included additional metrics like F1-score, recall, and precision to provide a more comprehensive evaluation, as relying solely on accuracy can be misleading. These metrics help assess the model’s performance in imbalanced scenarios and its ability to balance precision and recall.
TABLE IV: Average Models Metrics
Models Accuracy(%) F1 Score(%) Recall(%) Precision(%)
IBIS+SVM 98.95 98.78 98.72 98.85
Inception 88.12 88.01 87.49 91.52
Note: SVM - Support Vector Machine
Our method also showed a significant improvement over a transfer learning approach using a hybrid CNN-ABLSTM model from a related study[8]. Our proposed IBIS + SVM achieved an accuracy of 98.95% compared to their 85.82%, a remarkable 18.62% improvement under the same condition of 100 training epochs. This is because our approach uses a combination of signal processing (the SHARP algorithm for noise elimination), deep learning (the enhanced hybrid network for feature and temporal learning), and traditional machine learning (the SVM for refining classification). This pipeline and the inclusion of more attention layers, increased architectural complexity, and Doppler traces features, significantly enhanced generalization compared to the other study, which used raw unprocessed data.
IV-F Computational Complexity
A primary challenge with neural networks is computational cost, as a deeper network with more layers typically consumes more resources. To objectively measure this, Figure 8 shows the computational cost over 100 training epochs without using callbacks like early stopping.
Figure 8: Graph illustrating the execution time per training epoch for both the original Inception architecture and the proposed hybrid IBIS model.
We evaluated two distinct architectures:
  • The original Inception network;
  • An Inception network combined with a Bi-directional Long Short-Term Memory (BiLSTM) layer.
The standalone Inception network had the lowest computational cost and the shortest training time per epoch due to its shallow architecture. However, it failed to accurately classify movements, producing false positives and showing temporal instability.
V Conclusion
This study successfully addresses the challenge of Human Activity Recognition (HAR) using Wi-Fi sensing, a non-intrusive and low-cost technology with broad applications in healthcare, surveillance, and monitoring. We introduced a novel hybrid neural network architecture, IBIS, that combines Inception-BiLSTM with a SVM to enhance data generalization and classification accuracy. Our model’s performance was rigorously evaluated across scenarios with five and eight activities, achieving an exceptional average accuracy of nearly 99%. This not only demonstrates the method’s robustness against noise and interference but also represents a significant improvement of approximately 10.94% over the state-of-the-art SHARP study. The excellent results were further validated by high-performing ROC curves and other metrics.
We also conducted a comparative analysis, confirming that the Inception-BiLSTM + SVM combination outperformed other configurations. This underscores the effectiveness of our approach, which strategically leverages convolutional layers for feature extraction, BiLSTM for temporal analysis, and SVM for refining decision boundaries. For future work, we propose exploring refinements in the Doppler curves acquisition process, so that the complexity of the neural network training/post-processing can be reduced.